fr + h)-f(x) Find h for the given function. f(x) = 3x2-x+1 O 6x + 3h - 1 Oh o 3h2 +6hx - h O1 O 3 - 1
Answer #1

Solution:- We have, f() = 326?-X+1 11 f(x+h) = 3600+h)-(3+) + 1 = 3(32+2=cht h?) - be-h +1 - 30c2 + 6xht 3h²-2-htl - (38?_X +